# FluoCLIP: Stain-Aware Focus Quality Assessment in Fluorescence Microscopy

### v1

@@ 6.1 Datasets and Experiment Settings
Datasets. All experiments were conducted with consistent training protocols across FocusPath [11], BBBC006 [14], and FluoMix. Within FluoMix, brain tissue images were used for training, and images from other tissues were used for validation and few-shot testing to assess cross-tissue generalization.
Training Details. We adopted CLIP [16] as the vision-language backbone. Image encoder was initialized from either CLIP-pretrained ResNet50 [10] or ViT-B/16 [5] backbone, with its pretrained counterparts and random initialized prompts. Each baseline model was trained using the RAdam optimizer and a cosine learning rate scheduler. For FluoCLIP, we adopted a two-stage training protocol consisting of 10 epochs for the Stain-Grounding stage and 100 epochs for the Stain-Guided Ranking stage. All other training settings, including data preprocessing and augmentations, were kept identical across models. Input images were resized to , randomly cropped to , and augmented using horizontal flips and standard normalization.
@@ 6.2 Comparative Study
| Type | Configuration | Acc. (%) | Step Gain | Total Gain | |||||
| (A) | Baseline (OrdinalCLIP) with Rank () | ✓ | 83.12 | - | - | ||||
| (B) | (A) + Plain Stain Token () | ✓ | ✓ | 82.47 | -0.64 | - | |||
| (C) | (A) + Learnable Stain Token () | ✓ | ✓ | 80.05 | -2.62 | - | |||
| (D) | (A) + Grounded Stain Token () | ✓ | ✓ | 84.28 | 1.16 | 1.16 | |||
| (E) | (D) + Stain-Guided Rank Token () | ✓ | ✓ | 85.21 | 0.93 | 2.09 |
Results on FluoMix. FluoMix presents the greatest challenge due to its high stain variability and heterogeneous tissue structures. Traditional models exhibit significant drops in performance, with accuracy ranging from 47.261.99% to 63.461.38%, indicating their inability to adapt to diverse staining conditions. In contrast, FluoCLIP achieves 85.210.88% (Table 2), substantially outperforming all baselines. The expicit integration of stain-specific text tokens with stain-conditioned rank prompts allows FluoCLIP to model complex stain-dependent focus variations that conventional FQA models cannot capture. FluoCLIP also maintains its advantage with a ViT-B/16 encoder (Table 3), demonstrating strong robustness across architectures.
@@ 6.3 Ablation Study
Contribution of Individual Components. Explicit stain alignment and stain-conditioned ranking are critical to FluoCLIP. We validate this on FluoMix using a ResNet-50 CLIP backbone. Table 6 summarizes the results. The baseline (A), identical to OrdinalCLIP [13], achieves 83.12% accuracy. Adding a non-learnable stain token (B) slightly decreases accuracy by 0.64%, indicating that simply appending stain information is ineffective because CLIP has no prior knowledge of fluorescence-specific stain semantics. Even when the stain token is made learnable (C), accuracy drops more substantially (-2.62%), showing that introducing trainable text parameter without explicitly aligning them to visual features can inject uninformative or noisy semantics that interfere with learning. Introducing learnable stain tokens through the stain-grounding stage (D) improves accuracy to 84.28%, confirming the benefit of learning stain-visual correlations. Finally, incorporating the stain-guided ranking stage (E) raises accuracy to 85.21%, a total gain of +2.09% over the baseline. These results verify that both stages contribute synergistically—progressively enhancing the model’s ability to capture stain semantics and model stain-specific focus variations.
Effects of Learning Stain Representations. The ablations on BiomedCLIP [22] confirm the same trend observed in Table 6—explicitly understanding stains and modeling stain-guided ranks significantly improve performance (Table 7). Despite these gains, FluoCLIP with the standard CLIP ViT-B/16 backbone achieves a much higher accuracy of 90.75%, surpassing all BiomedCLIP variants. We attribute this gap to two limitations of BiomedCLIP: (1) text encoder (PubMedBERT), which lacks exposure to optical descriptors such as “sharp”, “blur” or actual non-optimal images, and (2) image encoder, which was trained on mixed biomedical imagery rather than fluorescence microscopy. These findings highlight the importance of explicitly learning stain-grounded textual embeddings, rather than relying on domain-specific pretraining alone.
@@ 7 Conclusion
We introduced the task of stain-aware ordinal focus quality assessment (FQA) for fluorescence microscopy, highlighting the importance of modeling stain-dependent focus behavior. To address this challenge, we proposed FluoCLIP, a two-stage vision–language framework that explicitly learns the interaction between fluorophore characteristics and focus quality. In the stain-grounding stage, learnable stain tokens are aligned with CLIP’s visual encoder, while the stain-guided ranking stage leverages these aligned embeddings for stain-conditioned ordinal reasoning. Together with our dataset FluoMix, which covers multi-stain, tissue-level fluorescence images, FluoCLIP establishes a new foundation for robust focus assessment under biologically diverse conditions. Extensive experiments on BBBC006 and FluoMix confirm that the two-stage learning scheme progressively enhances the model’s understanding of stain semantics and consistently outperforms conventional FQA methods and recent vision–language baselines.
While FluoCLIP demonstrates strong performance, it still relies on predefined stain categories during alignment, limiting adaptation to entirely unseen stains or imaging modalities. Future work will explore automatic stain discovery and cross-channel reasoning to further generalize stain-aware FQA across broader fluorescence domains.
@@ 8 Spatial Frequency
In Sec. 5, we used the spatial frequency (SF) [7] metric as a quantitative proxy for image sharpness to analyze stain-dependent focus behavior across datasets. For completeness, we provide the exact formulation of SF below.
@@ 9.1 Overview of Datasets
We evaluated FluoCLIP on three datasets used in Sec. 5 and Sec. 6: FocusPath [11], BBBC006 [14], and our fluorescence microscopy dataset FluoMix. To ensure consistent training across datasets, all z-stacks were reorganized into discrete focus levels, ranging from in-focus to severely defocused, following the ordinal labeling protocol described in the main paper (Sec. 4). Representative examples from each dataset are shown in Fig. 4 and Fig. 5. The former illustrates focus-level (rank) variations across datasets, whereas the latter highlights stain-dependent visual differences across the three datasets.
@@ 9.4 FluoMix
Each field of view is acquired as a full z-stack (34 planes for FluoMix) covering the progression from in-focus to severely defocused slices. For consistency with FocusPath and BBBC006, all z-stacks are converted into 10 ordinal focus levels (0 = in-focus, 9 = out-of-focus) using the relative labeling scheme described in Sec. 4. Representative examples are shown in Fig. 4.
@@
10 FluoMix Dataset: Biological Preparation and Imaging Protocols
@@ 11.3 Few-Shot Generalization to Unseen Stain & Tissue Pairs
We now evaluate the most challenging generalization setting, where both the stain and the tissue differ from those seen during base training. Starting from a checkpoint trained on the FluoMix brain-tissue subset (D1–D4), we adapt the model to two entirely unseen nuclear stains–Hoechst 33342 (D5–D6) and DAPI (D7)–each acquired from different tissues than the training domain. Adaptation uses only labeled samples per focus rank for each stain (), resulting in training images.
To ensure comparability with earlier experiments, we restrict evaluation to nucleus-targeting stains (DAPI, Hoechst 33342). Since these stains highlights the same biological structure, morphological differences caused by tissue variability are reduced, allowing stain-dependent optical differences to be examined more clearly (Fig. 5)
@@ 12 Ablation Study
We compare three types of stain tokens: is a non-learnable text token that simply inserts the stain name, is a learnable token updated during training without explicit grounding, and is our grounded stain token obtained through the two-stage procedure.
Table 13 varies two factors that jointly define the stain-token type—Fine-tune(S) (trainable vs. frozen tokens) and Two Stage (with vs. without stain grounding)—along with an additional orthogonal choice of whether to include the text-side Adapter after CLIP text encoder.
@@
Effects of Stain-Grounding. Configuration (A) uses non-learnable stain tokens (), while (C) replaces it with learnable stain tokens () but without the grounding stage (Two Stages). Interestingly, (C) performs worse than (A), indicating that simply making stain tokens trainable does not help. Without the grounding phase, the model receives no visual constraint linking stain-specific text features to fluorescence appearance, and learning tends to inject noise rather than meaningful stain semantics.
@@
Effects of Adapter. Configuration (B) and (D) evaluate the adapter with non-grounded stain tokens ( in (B), in (D)). The adapter yields moderate improvements in both cases, but the gains remain limited because the underlying stain tokens are not aligned with visual features. Thus, architectural augmentation alone cannot produce meaningful stain-aware embeddings—the adapter benefits only materialize when the stain tokens themselves are grounded.
Combined Effects. Configuration (E) and (F) corresponds to the two-stage FluoCLIP pipeline, with (F) additionally using the text-side adapter. Grounded stain tokens alone (E) already outperform all non-grounded alternatives (A, C), verifying that the grounding stage is the primary source of performance improvement. Adding the adapter (F) achieves the best results overall (84.28%), confirming that grounded stain semantics and the adapter contribute complementary benefits. Overall, these ablations demonstrate that (i) stain-grounding is necessary for producing stable and informative stain-aware embeddings, and (ii) the adapter further refines these representations, but works best when grounding is present. This explains why the full FluoCLIP pipeline achieves the strongest performance.
@@ 13 Algorithm
| Type | Configuration | Acc. (%) | |||||
| (A) | Baseline (OrdinalCLIP) with Rank () | ✓ | ✓ | 83.120.41 | |||
| (B) | (A) + Plain Stain Token () | ✓ | ✓ | 83.213.93 | |||
| (C) | (A) + Learnable Stain Token () | ✓ | ✓ | 81.380.63 | |||
| (D) | (A) + Grounded Stain Token () | ✓ | ✓ | 84.280.88 | |||
| (E) | (D) + Stain-Guided Rank Token | ✓ | ✓ | 85.210.88 |

### v2

@@ 6.1 Datasets and Experiment Settings
Datasets. All experiments were conducted with consistent training protocols across FocusPath [13], BBBC006 [17], and FluoMix. Within FluoMix, brain tissue images were used for training, and images from other tissues were used for validation and few-shot testing to assess cross-tissue generalization. More details are provided in the supplementary material.
Training Details. We adopted CLIP [19] as backbone, using either CLIP-pretrained ResNet50 [12] or ViT-B/16 [5] image encoders with pretrained weights. FluoCLIP is trained with a two-stage training protocol consisting of 10 epochs for the Stain-Grounding and 100 epochs for the Stain-Guided Ranking. All models are trained under identical settings for fair comparison, including data preprocessing and augmentations. Input images were resized to , randomly cropped to , and augmented using horizontal flips and standard normalization.
@@
| Type | Configuration | Acc. (%) | Step Gain | Total Gain | |||||
| (A) | Baseline (OrdinalCLIP) with Rank () | ✓ | 83.120.41 | - | - | ||||
| (B) | (A) + Plain Stain Token () | ✓ | ✓ | 83.213.93 | 0.09 | - | |||
| (C) | (A) + Learnable Stain Token () | ✓ | ✓ | 81.380.63 | -1.74 | - | |||
| (D) | (A) + Grounded Stain Token () | ✓ | ✓ | 84.280.88 | 1.16 | 1.16 | |||
| (E) | (D) + Stain-Guided Rank Token () | ✓ | ✓ | 85.210.88 | 0.93 | 2.09 |
@@ 6.2 Comparative Study
Results on FluoMix. FluoMix presents the greatest challenge due to its high stain variability and heterogeneous tissue structures. Traditional models exhibit significant drops in performance, with accuracy ranging from 47.261.99% to 63.461.38%, indicating their inability to adapt to diverse staining conditions. This suggests that purely visual models struggle to disentangle stain-induced variation from defocus, motivating the need for semantic conditioning. In contrast, FluoCLIP achieves 85.210.88% (Table 2), substantially outperforming all baselines. The explicit integration of stain-specific text tokens with stain-conditioned rank prompts allows FluoCLIP to model complex stain-dependent focus variations that conventional FQA models cannot capture. FluoCLIP also maintains its advantage with a ViT-B/16 encoder (Table 3), demonstrating strong robustness across architectures.
@@ 6.3 Ablation Study
Contribution of Individual Components. Explicit stain alignment and stain-conditioned ranking are critical to FluoCLIP. We validate this on FluoMix using a ResNet-50 CLIP backbone. Table 6 summarizes the results. The baseline (A), identical to OrdinalCLIP [15], achieves 83.12% accuracy. Adding a non-learnable stain token (B) slightly increases accuracy by 0.09%, indicating that simply appending stain information is ineffective because CLIP has no prior knowledge of fluorescence-specific stain semantics. Even when the stain token is made learnable (C), accuracy drops more substantially (-1.74%), showing that introducing trainable text parameter without explicitly aligning them to visual features can inject uninformative or noisy semantics that interfere with learning. Introducing learnable stain tokens through the stain-grounding stage (D) improves accuracy to 84.28%, confirming the benefit of learning stain-visual correlations. Finally, incorporating the stain-guided ranking stage (E) raises accuracy to 85.21%, a total gain of +2.09% over the baseline. These results verify that both stages contribute synergistically—progressively enhancing the model’s ability to capture stain semantics and model stain-specific focus variations.
Effects of Learning Stain Representations. The ablations on BiomedCLIP [25] confirm the same trend observed in Table 6—explicitly understanding stains and modeling stain-guided ranks significantly improve performance (Table 7). Despite these gains, FluoCLIP with the standard CLIP ViT-B/16 backbone achieves a much higher accuracy of 90.75%, surpassing all BiomedCLIP variants. We attribute this gap to two limitations of BiomedCLIP: (1) text encoder (PubMedBERT [10]), which lacks exposure to optical descriptors such as “sharp”, “blur” or actual non-optimal images, and (2) image encoder, which was trained on mixed biomedical imagery rather than fluorescence microscopy. These findings highlight the importance of explicitly learning stain-grounded textual embeddings, rather than relying on domain-specific pretraining alone.
@@ 7 Conclusion
We introduced stain-aware ordinal focus quality assessment (FQA) for fluorescence microscopy, demonstrating that focus behavior is inherently stain-dependent rather than governed by a single global ordering. To address this, we proposed FluoCLIP, a two-stage vision–language framework that learns stain grounding and enables stain-conditioned ordinal reasoning. Together with our dataset FluoMix, this establishes a practical foundation for modeling focus under biologically diverse fluorescence conditions. Extensive experiments show that FluoCLIP consistently outperforms both conventional FQA methods and recent vision–language baselines, highlighting the importance of explicitly modeling stain-dependent focus behavior rather than relying on shared ordinal structures. While FluoCLIP assumes predefined stain categories, future work will explore automatic stain discovery and more general conditioning mechanisms to extend stain-aware FQA to broader and less structured imaging settings.
Acknowledgment
This research was supported by the Bio & Medical Technology Development Program of the National Research Foundation (NRF) Korea (MSIT; RS-2022-NR068424), the G-LAMP Program of the NRF Korea (RS-2025-25442252), and the Korea Basic Science Institute (National Research Facilities and Equipment Center) funded by the Ministry of Education, Republic of Korea (RS-2024-00436263, RS-2025-02310437).
@@ 8 Spatial Frequency
In Section 5, we used the spatial frequency (SF) [7] metric as a quantitative proxy for image sharpness to analyze stain-dependent focus behavior across datasets. For completeness, we provide the exact formulation of SF below.
@@ 9.1 Overview of Datasets
We evaluated FluoCLIP on three datasets used in Section 5 and Section 6: FocusPath [13], BBBC006 [17], and our fluorescence microscopy dataset FluoMix. To ensure consistent training across datasets, all z-stacks were reorganized into discrete focus levels, ranging from in-focus to severely defocused, following the ordinal labeling protocol described in the main paper (Section 4). Representative examples from each dataset are shown in Figure 5 and Figure 6. The former illustrates focus-level (rank) variations across datasets, whereas the latter highlights stain-dependent visual differences across the three datasets.
@@ 9.4 FluoMix
Each field of view is acquired as a full z-stack (34 planes for FluoMix) covering the progression from in-focus to severely defocused slices. For consistency with FocusPath and BBBC006, all z-stacks are converted into 10 ordinal focus levels (0 = in-focus, 9 = out-of-focus) using the relative labeling scheme described in Section 4. Representative examples are shown in Figure 5.
@@
10 Biological Preparation / Imaging Protocols
@@ 11.3 Few-Shot Generalization to Unseen Stain & Tissue Pairs
We now evaluate the most challenging generalization setting, where both the stain and the tissue differ from those seen during base training. Starting from a checkpoint trained on the FluoMix brain-tissue subset (D1–D4), we adapt the model to two entirely unseen nuclear stains–Hoechst 33342 (D5–D6) and DAPI (D7)–each acquired from different tissues than the training domain. Adaptation uses only labeled samples per focus rank for each stain (), resulting in training images. To ensure comparability with earlier experiments, we restrict evaluation to nucleus-targeting stains (DAPI, Hoechst 33342). Since these stains highlights the same biological structure, morphological differences caused by tissue variability are reduced, allowing stain-dependent optical differences to be examined more clearly (Figure 6)
@@ 12 Ablation Study
We compare three types of stain tokens: is a non-learnable text token that simply inserts the stain name, is a learnable token updated during training without explicit grounding, and is our grounded stain token obtained through the two-stage procedure. Table 13 varies two factors that jointly define the stain-token type—Fine-tune () (trainable vs. frozen tokens) and Two Stage (with vs. without stain grounding)—along with an additional orthogonal choice of whether to include the text-side Adapter after CLIP text encoder.
@@
12.1 Effects of Stain-Grounding
Configuration (A) uses non-learnable stain tokens (), while (C) replaces it with learnable stain tokens () but without the grounding stage (Two Stages). Interestingly, (C) performs worse than (A), indicating that simply making stain tokens trainable does not help. Without the grounding phase, the model receives no visual constraint linking stain-specific text features to fluorescence appearance, and learning tends to inject noise rather than meaningful stain semantics.
@@ 12.1 Effects of Stain-Grounding
12.2 Effects of Adapter
Configuration (B) and (D) evaluate the adapter with non-grounded stain tokens ( in (B), in (D)). The adapter yields moderate improvements in both cases, but the gains remain limited because the underlying stain tokens are not aligned with visual features. Thus, architectural augmentation alone cannot produce meaningful stain-aware embeddings—the adapter benefits only materialize when the stain tokens themselves are grounded.
12.3 Combined Effects
Configuration (E) and (F) corresponds to the two-stage FluoCLIP pipeline, with (F) additionally using the text-side adapter. Grounded stain tokens alone (E) already outperform all non-grounded alternatives (A, C), verifying that the grounding stage is the primary source of performance improvement. Adding the adapter (F) achieves the best results overall (84.28%), confirming that grounded stain semantics and the adapter contribute complementary benefits. Overall, these ablations demonstrate that (i) stain-grounding is necessary for producing stable and informative stain-aware embeddings, and (ii) the adapter further refines these representations, but works best when grounding is present. This explains why the full FluoCLIP pipeline achieves the strongest performance.
@@ 13 Algorithm
14 Biomedical CLIP Variants
To evaluate whether domain-specific pretraining alone suffices for fluorescence FQA, we compare FluoCLIP with biomedical CLIP variants, including PubMedCLIP [8] and PMC-CLIP [16]. These models are pretrained on large-scale biomedical image–text corpora and provide stronger domain alignment than the original CLIP.
As shown in Table 14, both PubMedCLIP and PMC-CLIP achieve competitive performance on FluoMix, outperforming standard image-only baselines. In particular, PMC-CLIP benefits from broader biomedical coverage and achieves higher accuracy than PubMedCLIP.
However, FluoCLIP consistently outperforms both models, achieving the highest accuracy. This result indicates that domain-specific pretraining alone is insufficient for FQA, as it does not explicitly capture stain-dependent optical variations or ordinal focus structure. By incorporating stain-aware grounding and stain-conditioned ordinal modeling, FluoCLIP effectively resolves these limitations and yields more consistent and accurate focus predictions across diverse fluorescence conditions.
| Method | PubMedCLIP | PMC-CLIP | FluoCLIP |
| Accuracy (%) | 81.480.34 | 84.060.23 | 85.210.88 |